Rica and Jasmin converted a clock to become a spinner, and then they decided to play a game. If Jasmin decided to choose an even number or a number greater than 8, what is the probability that she will win the game?


1
Expert's answer
2021-05-17T01:41:02-0400

If she choose the even number, then the probability for her to win is 6/12=0.5

If she choose a number greater than 8 (which is 9, 10, 11 or 12, total 4 out of 12), then the probability is 4/12=0.33

If she choose both at the same time, than (2, 4, 6, 8, 9, 10, 11, 12, total 8 out of 12), then the probability is 8/12=0.67